### v2

@@ 2.2 Writing Examples
Why do companies in the United Kingdom pay scrip (i.e., in stock) dividends? Are tax savings the sole motive for this option? If so, are all tax-loss companies offering this option? Or is this option driven by other motives, such as cash savings, signaling, and agency conflicts? In this study, I provide some insights into the motivation for scrip-dividend payment by comparing the operational performance and other characteristics of all companies that distributed scrip dividends with those of a control group of non-scrip-paying, but otherwise similar, firms. Scrip dividends are offered by an increasing number of companies in the United Kingdom as an option whereby shareholders are able to choose between receiving dividends in cash or the equivalent in the form of shares (scrip). Companies stress tax savings when they offer this option because, unlike cash dividends, the scrip option is not subject to a payment of the advanced corporation tax (ACT). I find that there are no significant differences in the scrip and non-scrip-paying firms’ tax exposures. Compared to the control sample, firms that issue scrip dividends do not report a higher proportion of irrecoverable ACT in their accounts, nor do they generate a high proportion of their earnings overseas, earnings on which the ACT cannot be claimed. My results also show that this option is not driven by cash shortages, nor is it a substitute for external finance and/or a cut in cash dividends. However, I find that firms that pay scrip dividends are, on average, large, and have high dividend yield, suggesting that the cash saved is substantial. However, these firms already have high cash-flow balances and low growth opportunities. The overall results suggest that firms in the UK do not appear to be making the optimal financial choice for their dividend policies. Like stock dividends and dividend reinvestment plans in the US. scrip dividends provide both firm and shareholders with many benefits. Companies are able to retain cash without altering their payout policies or raising new funds in the capital markets; thus, companies that use this option can save on borrowing costs, underwriting fees, other issue costs, and avoid negative signals of new equity. At the same time, they provide their shareholders with the opportunity to increase their holdings without incurring any transaction costs. In addition, the UK institutional setting and relevant tax code provisions allow companies to derive tax benefits from this option. Unlike cash dividends, this option is not subject to ACT. This tax is first paid when a company declares cash dividends and is deducted from its corporation-tax liability nine months after the accounting year-end if dividends are paid from domestic earnings and if taxable profit is higher than gross cash dividends. If these two conditions are not met, surplus ACT is carried in the accounts and set off against preceding or immediately following periods. For companies with no previous or foreseeable future taxable profit, ACT is written off as a loss against reserves. Thus, companies with prospective ACT loss are more likely to pay scrip, rather than cash, dividends. However, scrip dividends are not costless. Firms incur administrative costs in Petting up and running the scheme, such as the costs of advertising the option, preparing and mailing the scrip-dividend prospectus, and printing and allocating the new shares. Moreover, scrip dividends impose costs on the firm’s shareholders. This option is not offered to foreign investors. For some domestic investors, it is taxed differently from cash dividends. While the tax credit associated with cash dividends can be claimed by all investors, the tax credit on scrip dividends can only be claimed by tax-paying individual investors. Individuals who have reliefs and allowances in excess of their income, as well as corporate investors, forgo the tax credit when they opt for scrip dividends. Thus, corporate and nontaxable individual investors are likely to prefer cash dividends; individual shareholders taxed at a higher rate of income tax may opt for scrip dividends for which the firm issues additional shares. However, such additional shares could result in a dilution of control for shareholders who opt for cash dividends, and a loss in future dividend increases if scrip dividends limit the scope for such increases. Why, then, do companies issue scrip dividends? One possibility is that individual tax-paying investors request this option from their companies, which suggests that companies are subject to monitoring by atomistic shareholders while large shareholders, such as corporate investors, are passive owners. I suggest a number of areas for further research that will assess the extent to which the granting of this option is driven by nonfinancial considerations. The rich variety of securities innovations in recent years continues to intrigue academicians and practitioners alike. The array of innovative securities includes a variety of equity-linked debt securities. On January 28, 1991, the Republic of Austria publicly offered \$100 million principal amount of a new equity-linked debt security called stock index growth notes (”SIGNs”) in the United States. The SIGNs were scheduled to mature in approximately 5.5 years. They would make no payments of interest prior to maturity. If the value of the S&P 500 is below 336.69 on the maturity date, the holder will receive \$10, the initial offering price. If the value exceeds 336.69, the holder will get \$10 plus \$10 multiplied by the percentage appreciation in the S&P 500 above 336.69. SIGNS may thus be characterized as a package consisting of (i) a 5.5-year triple-A-rated zero coupon note plus (ii) a 5.5-year European call option, or warrant, on the S&P 500 with a strike price of 336.69. The value of each SIGN can be expressed as the sum of (i) the value of the zero coupon bond component plus (ii) the value of the call option component plus (iii) the value resulting from interest on the SIGNs not being taxable until the maturity date plus (iv) the reduction, if any, in transaction costs vis-a-vis acquiring a comparable zero coupon bond and purchasing call options on the S&P 500 separately plus (v) the value, if any, attributable to creating an investment alternative that is not otherwise available to the investors who purchased the SIGNs. On January 28. 1991, the Republic of Austria had outstanding an issue of zero coupon bonds maturing July 17, 1995, that was yielding 8.05% per annum semiannually compounded. Based on this yield, the zero coupon bond component of each SIGN was worth approximately \$6.48. The call option component, valued using the Black-Scholes model modified to account for dividend-paying stocks, was estimated to be \$2.30. The value of the tax arbitrage created through the introduction of SIGNs equals the sum of (i) the tax benefit resulting from the investor’s ability to defer income taxes on the amortization of the zero coupon bond component (\$0.11 per SIGN) plus (ii) the tax benefit resulting from the investor’s ability to defer income taxes that would otherwise be incurred on the option component under the mark-to-market rules (\$0.09 per SIGN). The total value of the tax arbitrage is \$0.20 per SIGN, just two percent of the public offering price per SIGN. I compared the relative transaction costs associated with (i) purchasing 1,000 SIGNs (\$0.02 per SIGN), and (ii) purchasing a Treasury strip with \$10,000 principal amount and a portfolio that replicates the call option component of each SIGN (\$0.16 per SIGN). The present-value after-tax transaction cost savings are thus roughly \$0.14 per SIGN. The creation of a new security can benefit investors if it creates new risk-return combinations. Call options on the S&P 500 with a time to expiration of 5.5 years are not available on any of the options exchanges. An investor who wished to duplicate the return stream of the SIGNs could try to purchase a customized option in the over-the-counter market or employ a dynamic replication strategy. The SIGNs were sold in small denominations primarily to retail investors, who would not have had access to the over-the-counter market on account of the small transaction size. Dynamic replication would be prohibitively expensive for at least some retail investors because of transaction costs. The SIGNs were sold to investors at a price of \$10.00. Based on the four estimated component values. the implied value attributable to creating a new investment alternative is \$0.88 per SIGN. However, this value also includes the effect of any mispricing that might have occurred. The predicted value of the SIGNs was 8.80% below the market price on January 28, 1991. At the end of February 1991, the predicted value was only 1.60% below the market price. At the end of March 1991, the predicted value was \$10.03, which slightly exceeded the \$10,00 market price. Thereafter, the predicted value and the market price approximate one another. This price behavior suggests that investors may have initially overvalued the SIGNs but that the mispricing was eliminated within roughly one to two months. It also suggests that the value of the innovation should be attributed primarily to the value of the tax arbitrage and to the reduction in after-tax transaction costs it made possible. The Republic of Austria used approximately 30% of the \$9.50 net proceeds from the sale of the SIGNs issue to hedge its contingent S&P 500 liability, leaving proceeds net of hedging costs amounting to \$6.65. These net proceeds exceed the \$6.48 value of the bond component of each SIGN. The profit of \$0.17 per SIGN effectively reduced the Republic of Austria’s cost of issuing 5.5-year zero coupon debt to 7.56%, resulting in a savings of 49 basis points per annum. FemaleMaleInvolvedInformational
@@ 2.3 Writing Analysis
Table 1 shows the breakdown of informational and involved features for academic papers and patents, respectively. The results there suggest that, in absolute terms, informational features are used significantly more often in both text sources, by both men and women. This distribution seems plausible given the nature of our corpus (i.e., scientific articles and patents are generally concerned with communicating information), and is also consistent with prior work; Biber, (1988), for example, situated academic prose closer to the informational dimension, and Argamon et al. (2013) noted the “greater quantification inherent to most non-fiction genres.” Table 2 also shows the mean and standard deviation of involved and informational features combined and the Involved-Informational Ratio in the academic papers and patents samples.
@@
While the patterns we observe appear to hold universally across fields of science and technology, we find heterogeneity in the magnitude of the effect. Figure S9 in the Supplementary Materials shows the predicted gender differences by field. The differences are greater in the Social Sciences and Arts & Humanities and less pronounced in Physical Sciences and Technology, which may be explained by the fact that the former fields generally permit greater freedom of stylistic expression and higher use of involved features than the latter fields, which have more restricted stylistic rules (Argamon et al.,, 2003). Full regression results for each field are presented in Table LABEL:detailed_papers (for academic papers) and in Table LABEL:detailed_patents (for patents) in the Supplementary Materials.
@@ 2.4 Citation Analysis
Next, we turn to our analysis of gender differences in citations, the results of which are shown in Tables 5 (academic papers) and 6 (patents). We also provide a visualization of our findings in Figure 4. To simplify the presentation, in this section, we only report models predicting citation as a function of the rate of informational and involved feature utilization (i.e., we do not report the results of models that use the Involved-Informational Ratio as a predictor). We found substantially similar results, however, using the ratio measure.
@@
Overall, we observe significant gender homophily in citation patterns. Papers written by female authors receive a greater share of citations from papers with a female first author ($\beta=4.33,p<0.001$) and papers with a female last author ($\beta=3.57;p<0.001$). The opposite pattern holds for sample papers written by men.
@@
Taking into account the linguistic features, we see a consistent and statistically significant pattern across the outcome measures: papers with higher rates of utilization of involved features in their abstracts tend to be cited at a higher rate by papers with a female first author ($\beta=0.25,p<0.001$, Model 1) and papers with a female last author ($\beta=0.29,p<0.001$, Model 3). More substantively, we can interpret the coefficient estimates as indicating that for each unit increase in the rate of involved features used, the rate of citation by papers with a female first and last author is predicted to increase by 0.25 and 0.29, respectively. When we consider the use of informational features, however, the pattern flips ($\beta=-0.08,p<0.001$, Model 1 for citation by papers with a female first author; $\beta=-0.06,p<0.001$, Model 3 for citation by papers with a female first author). Papers with higher rates of utilization of involved features are cited at a lower rate by papers with a male first author ($\beta=-0.18,p<0.001$, Model 2) and papers with a male last author ($\beta=-0.30,p<0.001$, Model 4). We find that the opposite holds for utilization of informational features, although only in models predicting citation by papers with a male first author ($\beta=0.03,p<0.01$, Model 2).
@@
Turning to patents, we find results that are generally similar, but much weaker. While patents that use more informational features are less likely to be cited by patents with a female first ($\beta=-0.07,p<0.001$, Model 1) or last ($\beta=-0.07,p<0.001$, Model 3) inventor, none of the remaining coefficients tracking the utilization of informational or involved features are statistically significant. These results are not surprising, considering that patent descriptions, as a type of text, allow much less stylistic expression than academic papers do.
@@
Figure 4: Regression Coefficients for Citation Patterns. This panel displays plots that illustrate the magnitude and direction of the regression coefficients when the outcome variable is ‘Female first author’, ‘Female second author’, ‘Male first author’, or ‘Male second author’. (A) This plot reveals that papers written by female authors receive a greater share of citations from papers with a female first author and papers with a female last author; the opposite pattern holds for papers written by men. (B) This plot indicates that papers with higher rates of involved features tend to be cited at a higher rate by papers with a female first author and papers with a female last author; whereas papers with higher rates of involved features are cited at a lower rate by papers with a male first author. The involved rate was an insignificant predictor across all models for patents. (C) This plot shows that papers and patents authored by females, as either the first or last author, have a lower rate of informational features. Papers with a male first author were more likely to utilize informational features. But, informational rates are statistically insignificant for patents authored by males both as first and last authors. Note: The regression coefficients on the y-axis display a wider range for (A) compared to (B) and (C). The stars above each bar denote levels of significance (${*}p<0.05;{**}p<0.01;{***}p<0.001$), and ‘n.s.’ indicates ‘not significant’.
Table 5: Involved and informational rates for female and male authors by citation patterns. Regression results for predicted outcomes if author is female, with the following dependent variables: rate of citations from papers with a female, male first author and with a female, male last author. These results represent data across all fields.
(1) (2) (3) (4)
DV: Rate of citations from papers with a female first author DV: Rate of citations from papers with a male first author DV: Rate of citations from papers with a female last author DV: Rate of citations from papers with a male last author
@@
Author is female (1 = Yes) 4.33*** -4.33*** 3.57*** -3.64***
(0.07) (0.08) (0.07) (0.08)
Paper is cited 0 times (1 = Yes) -29.11*** -51.19*** -25.80*** -55.35***
(0.08) (0.08) (0.08) (0.08)
Constant 27.16*** 53.38*** 23.12*** 59.00***
(0.20) (0.22) (0.19) (0.22)
Field fixed effects Yes Yes Yes Yes
Year fixed effects Yes Yes Yes Yes
N 512520 512520 512520 512520
r2 0.3084 0.4720 0.2748 0.5244
  • Notes: Robust standard errors are shown in parentheses; p-values correspond to two-tailed tests.

### v4

@@ 3.2 Writing Examples
Why do companies in the United Kingdom pay scrip (i.e., in stock) dividends? Are tax savings the sole motive for this option? If so, are all tax-loss companies offering this option? Or is this option driven by other motives, such as cash savings, signaling, and agency conflicts? In this study, I provide some insights into the motivation for scrip-dividend payment by comparing the operational performance and other characteristics of all companies that distributed scrip dividends with those of a control group of non-scrip-paying, but otherwise similar, firms. Scrip dividends are offered by an increasing number of companies in the United Kingdom as an option whereby shareholders are able to choose between receiving dividends in cash or the equivalent in the form of shares (scrip). Companies stress tax savings when they offer this option because, unlike cash dividends, the scrip option is not subject to a payment of the advanced corporation tax (ACT). I find that there are no significant differences in the scrip and non-scrip-paying firms’ tax exposures. Compared to the control sample, firms that issue scrip dividends do not report a higher proportion of irrecoverable ACT in their accounts, nor do they generate a high proportion of their earnings overseas, earnings on which the ACT cannot be claimed. My results also show that this option is not driven by cash shortages, nor is it a substitute for external finance and/or a cut in cash dividends. However, I find that firms that pay scrip dividends are, on average, large, and have high dividend yield, suggesting that the cash saved is substantial. However, these firms already have high cash-flow balances and low growth opportunities. The overall results suggest that firms in the UK do not appear to be making the optimal financial choice for their dividend policies. Like stock dividends and dividend reinvestment plans in the US. scrip dividends provide both firm and shareholders with many benefits. Companies are able to retain cash without altering their payout policies or raising new funds in the capital markets; thus, companies that use this option can save on borrowing costs, underwriting fees, other issue costs, and avoid negative signals of new equity. At the same time, they provide their shareholders with the opportunity to increase their holdings without incurring any transaction costs. In addition, the UK institutional setting and relevant tax code provisions allow companies to derive tax benefits from this option. Unlike cash dividends, this option is not subject to ACT. This tax is first paid when a company declares cash dividends and is deducted from its corporation-tax liability nine months after the accounting year-end if dividends are paid from domestic earnings and if taxable profit is higher than gross cash dividends. If these two conditions are not met, surplus ACT is carried in the accounts and set off against preceding or immediately following periods. For companies with no previous or foreseeable future taxable profit, ACT is written off as a loss against reserves. Thus, companies with prospective ACT loss are more likely to pay scrip, rather than cash, dividends. However, scrip dividends are not costless. Firms incur administrative costs in Petting up and running the scheme, such as the costs of advertising the option, preparing and mailing the scrip-dividend prospectus, and printing and allocating the new shares. Moreover, scrip dividends impose costs on the firm’s shareholders. This option is not offered to foreign investors. For some domestic investors, it is taxed differently from cash dividends. While the tax credit associated with cash dividends can be claimed by all investors, the tax credit on scrip dividends can only be claimed by tax-paying individual investors. Individuals who have reliefs and allowances in excess of their income, as well as corporate investors, forgo the tax credit when they opt for scrip dividends. Thus, corporate and nontaxable individual investors are likely to prefer cash dividends; individual shareholders taxed at a higher rate of income tax may opt for scrip dividends for which the firm issues additional shares. However, such additional shares could result in a dilution of control for shareholders who opt for cash dividends, and a loss in future dividend increases if scrip dividends limit the scope for such increases. Why, then, do companies issue scrip dividends? One possibility is that individual tax-paying investors request this option from their companies, which suggests that companies are subject to monitoring by atomistic shareholders while large shareholders, such as corporate investors, are passive owners. I suggest a number of areas for further research that will assess the extent to which the granting of this option is driven by nonfinancial considerations. The rich variety of securities innovations in recent years continues to intrigue academicians and practitioners alike. The array of innovative securities includes a variety of equity-linked debt securities. On January 28, 1991, the Republic of Austria publicly offered \$100 million principal amount of a new equity-linked debt security called stock index growth notes (”SIGNs”) in the United States. The SIGNs were scheduled to mature in approximately 5.5 years. They would make no payments of interest prior to maturity. If the value of the S&P 500 is below 336.69 on the maturity date, the holder will receive \$10, the initial offering price. If the value exceeds 336.69, the holder will get \$10 plus \$10 multiplied by the percentage appreciation in the S&P 500 above 336.69. SIGNS may thus be characterized as a package consisting of (i) a 5.5-year triple-A-rated zero coupon note plus (ii) a 5.5-year European call option, or warrant, on the S&P 500 with a strike price of 336.69. The value of each SIGN can be expressed as the sum of (i) the value of the zero coupon bond component plus (ii) the value of the call option component plus (iii) the value resulting from interest on the SIGNs not being taxable until the maturity date plus (iv) the reduction, if any, in transaction costs vis-a-vis acquiring a comparable zero coupon bond and purchasing call options on the S&P 500 separately plus (v) the value, if any, attributable to creating an investment alternative that is not otherwise available to the investors who purchased the SIGNs. On January 28. 1991, the Republic of Austria had outstanding an issue of zero coupon bonds maturing July 17, 1995, that was yielding 8.05% per annum semiannually compounded. Based on this yield, the zero coupon bond component of each SIGN was worth approximately \$6.48. The call option component, valued using the Black-Scholes model modified to account for dividend-paying stocks, was estimated to be \$2.30. The value of the tax arbitrage created through the introduction of SIGNs equals the sum of (i) the tax benefit resulting from the investor’s ability to defer income taxes on the amortization of the zero coupon bond component (\$0.11 per SIGN) plus (ii) the tax benefit resulting from the investor’s ability to defer income taxes that would otherwise be incurred on the option component under the mark-to-market rules (\$0.09 per SIGN). The total value of the tax arbitrage is \$0.20 per SIGN, just two percent of the public offering price per SIGN. I compared the relative transaction costs associated with (i) purchasing 1,000 SIGNs (\$0.02 per SIGN), and (ii) purchasing a Treasury strip with \$10,000 principal amount and a portfolio that replicates the call option component of each SIGN (\$0.16 per SIGN). The present-value after-tax transaction cost savings are thus roughly \$0.14 per SIGN. The creation of a new security can benefit investors if it creates new risk-return combinations. Call options on the S&P 500 with a time to expiration of 5.5 years are not available on any of the options exchanges. An investor who wished to duplicate the return stream of the SIGNs could try to purchase a customized option in the over-the-counter market or employ a dynamic replication strategy. The SIGNs were sold in small denominations primarily to retail investors, who would not have had access to the over-the-counter market on account of the small transaction size. Dynamic replication would be prohibitively expensive for at least some retail investors because of transaction costs. The SIGNs were sold to investors at a price of \$10.00. Based on the four estimated component values. the implied value attributable to creating a new investment alternative is \$0.88 per SIGN. However, this value also includes the effect of any mispricing that might have occurred. The predicted value of the SIGNs was 8.80% below the market price on January 28, 1991. At the end of February 1991, the predicted value was only 1.60% below the market price. At the end of March 1991, the predicted value was \$10.03, which slightly exceeded the \$10,$0$ market price. Thereafter, the predicted value and the market price approximate one another. This price behavior suggests that investors may have initially overvalued the SIGNs but that the mispricing was eliminated within roughly one to two months. It also suggests that the value of the innovation should be attributed primarily to the value of the tax arbitrage and to the reduction in after-tax transaction costs it made possible. The Republic of Austria used approximately 30% of the \$9.50 net proceeds from the sale of the SIGNs issue to hedge its contingent S&P 500 liability, leaving proceeds net of hedging costs amounting to \$6.65. These net proceeds exceed the \$6.48 value of the bond component of each SIGN. The profit of \$0.17 per SIGN effectively reduced the Republic of Austria’s cost of issuing 5.5-year zero coupon debt to 7.56%, resulting in a savings of 49 basis points per annum. FemaleMaleInvolvedInformational
@@ 3.3 Writing Analysis
Table 1 shows the breakdown of informational and involved features for academic papers and patents, respectively. The results there suggest that, in absolute terms, informational features are used significantly more often in both text sources, by both men and women. This distribution seems plausible given the nature of our corpus (i.e., scientific articles and patents are generally concerned with communicating information), and is also consistent with prior work; biber1988, for example, situated academic prose closer to the informational dimension, and Argamon et al. (2013) noted the “greater quantification inherent to most non-fiction genres.” Table 2 also shows the mean and standard deviation of involved and informational features combined and the Involved-Informational Ratio in the academic papers and patents samples.
@@
While the patterns we observe appear to hold universally across fields of science and technology, we find heterogeneity in the magnitude of the effect. Figure LABEL:fig:Coefficients in the Supplementary Materials shows the predicted gender differences by field. The differences are greater in the Social Sciences and Arts & Humanities and less pronounced in Physical Sciences and Technology, which may be explained by the fact that the former fields generally permit greater freedom of stylistic expression and higher use of involved features than the latter fields, which have more restricted stylistic rules (argamon2003). Full regression results for each field are presented in Table LABEL:detailed_papers (for academic papers) and in Table LABEL:detailed_patents (for patents) in the Supplementary Materials.
@@ 3.4 Citation Analysis
Next, we turn to our analysis of gender differences in citations. To simplify the presentation, in this section, we only report models predicting citation as a function of the rate of informational and involved feature utilization (i.e., we do not report the results of models that use the Involved-Informational Ratio as a predictor). We found substantially similar results, however, using the ratio measure.
We present a set of nested regression models to assess whether linguistic variables mediate the relationship between author gender and citation counts. We compare a base model that includes author gender and relevant controls (field and year fixed effects, citation-zero flag) to a full model that adds our two key linguistic style measures. These results are shown in Table 3.4 (unstandardized coefficients for papers) and Table LABEL:table:RegressionsCitationsByGenderOfCitingInventorPatentsSeparateMeasuresV4Unstandardized (unstandardized coefficients for patents). See Supplementary Tables LABEL:table:RegressionsCitationsByGenderOfCitingAuthorPapersSeparateMeasuresV4Standardized and LABEL:table:RegressionsCitationsByGenderOfCitingInventorPatentsSeparateMeasuresV4Standardized for analyses with standardized coefficients. We also provide a visualization of our findings in Figure 5.
@@
Overall, we observe significant gender homophily in citation patterns. Papers written by female authors receive a greater share of citations from papers with a female first author ($\beta=4.40,p<0.001$, Model 1; $\beta=4.33,p<0.001$, Model 2) and papers with a female last author ($\beta=3.65;p<0.001$, Model 5; $\beta=3.57,p<0.001$, Model 6). The opposite pattern holds for sample papers written by men.
@@
Taking into account the linguistic features, we see a consistent and statistically significant pattern across the outcome measures: papers with higher rates of utilization of involved features in their abstracts tend to be cited at a higher rate by papers with a female first author ($\beta=0.25,p<0.001$, Model 2) and papers with a female last author ($\beta=0.29,p<0.001$, Model 6). More substantively, we can interpret the coefficient estimates as indicating that for each unit increase in the rate of involved features used, the rate of citation by papers with a female first and last author is predicted to increase by 0.25 and 0.29, respectively. When we consider the use of informational features, however, the pattern flips ($\beta=-0.08,p<0.001$, Model 2 for citation by papers with a female first author; $\beta=-0.06,p<0.001$, Model 6 for citation by papers with a female last author). Papers with higher rates of utilization of involved features are cited at a lower rate by papers with a male first author ($\beta=-0.18,p<0.001$, Model 4) and papers with a male last author ($\beta=-0.30,p<0.001$, Model 8). We find that the opposite holds for utilization of informational features, although only in models predicting citation by papers with a male first author ($\beta=0.03,p<0.01$, Model 4).
@@
Turning to patents, we find results that are generally similar, but much weaker. While patents that use more informational features are less likely to be cited by patents with a female first ($\beta=-0.07,p<0.001$, Model 2) or last ($\beta=-0.07,p<0.001$, Model 6) inventor, none of the remaining coefficients tracking the utilization of informational or involved features are statistically significant. These results are not surprising, considering that patent abstracts, as a type of text, allow much less stylistic expression than academic papers do.
Across all specifications, we find that the inclusion of linguistic variables leads to a small but consistent increase in explained variance — for instance, in the model predicting citations from female first authors, the $R^{2}$ increases from 0.3080 (Table 3.4, Model 1) to 0.3084 (Model 2). At the same time, we observe a modest reduction in the author gender coefficient: the effect of being female declines from 4.40 to 4.33, an attenuation of approximately 1.6%. Similar reductions appear across other citing-author-gender models—for example, a 2.19% attenuation for citations from female last authors (Models 5 vs. 6), and 1.89% for male last authors (Models 7 vs. 8). These changes, though modest, are consistent with a partial mediation pathway, suggesting that gendered language partially explains the observed citation disparities by author gender.
Although the increase in $R^{2}$ is small in absolute terms, the standardized coefficients on linguistic variables show that minor changes in writing style can have meaningful effects on citation outcomes. For example, in Model 2 (Table LABEL:table:RegressionsCitationsByGenderOfCitingAuthorPapersSeparateMeasuresV4Standardized), a one standard deviation increase in involved language is associated with a 0.52 standard deviation increase in citations from female first authors. Given that a one-unit change corresponds to just one additional involved word per 100 words, even modest stylistic shifts—for example, increasing involved language by a few words per 100—has significant implications. These effects, while subtle, suggest that stylistic presentation contributes to scholarly recognition, particularly along gendered lines.
Finally, we note that these patterns align with our theoretical expectations—male authors appear less likely to cite involved language and more likely to cite informational language, while female authors exhibit the opposite tendency. These findings suggest that linguistic style may function not only as a marker of epistemic stance, but also as a gendered cue that interacts with how work is received and cited. Taken together, the nested models offer greater transparency into the structure of these relationships and help clarify the mechanisms by which gendered citation disparities may arise.
Controlling for Content Similarity. We also examined whether our results hold when controlling for content similarity between cited and citing abstracts. To do this, we conducted additional regression analyses that included not only observed citations but also counterfactual citations—cases in which a paper or patent did not cite the paper or patent in question but was at risk of doing so due to high content similarity. In this modified analysis, we generally find that informational and involved rates continue to predict citation patterns, depending on whether the citing paper or patent has a female or male first or last author. A detailed account of these analyses is provided in the Controlling for Content Similarity section of the Supplementary Materials.
@@
Figure 5: Regression Coefficients for Citation Patterns. This panel displays plots that illustrate the magnitude and direction of the regression coefficients when the outcome variable is ‘Female first author’, ‘Female second author’, ‘Male first author’, or ‘Male second author’. (A) This plot reveals that papers written by female authors receive a greater share of citations from papers with a female first author and papers with a female last author; the opposite pattern holds for papers written by men. (B) This plot indicates that papers with higher rates of involved features tend to be cited at a higher rate by papers with a female first author and papers with a female last author; whereas papers with higher rates of involved features are cited at a lower rate by papers with a male first author. The involved rate was an insignificant predictor across all models for patents. (C) This plot shows that papers and patents authored by females, as either the first or last author, have a lower rate of informational features. Papers with a male first author were more likely to utilize informational features. But, informational rates are statistically insignificant for patents authored by males both as first and last authors. Note: The regression coefficients on the y-axis display a wider range for (A) compared to (B) and (C). The stars above each bar denote levels of significance (p<0.05;p<0.01;p<0.001{*}p<0.05;{**}p<0.01;{***}p<0.001), and ‘n.s.’ indicates ‘not significant’.
Citations from female first author Citations from male first author Citations from female last author Citations from male last author
(1) (2) (3) (4) (5) (6) (7) (8)
Author is female (1 = Yes) 4.40*** 4.33*** -4.38*** -4.33*** 3.65*** 3.57*** -3.71*** -3.64***
(0.07) (0.07) (0.08) (0.08) (0.07) (0.07) (0.08) (0.08)
Paper is cited 0 times (1 = Yes) -29.15*** -29.11*** -51.18*** -51.19*** -25.82*** -25.80*** -55.37*** -55.35***
(0.08) (0.08) (0.08) (0.08) (0.07) (0.08) (0.08) (0.08)
@@
Constant 26.98*** 27.16*** 52.99*** 53.38*** 23.55*** 23.12*** 57.42*** 59.00***
(0.07) (0.20) (0.08) (0.22) (0.07) (0.19) (0.08) (0.22)
Field fixed effects Yes Yes Yes Yes Yes Yes Yes Yes
Year fixed effects Yes Yes Yes Yes Yes Yes Yes Yes
N 512520 512520 512520 512520 512520 512520 512520 512520
R-squared 0.3080 0.3084 0.4718 0.4720 0.2743 0.2748 0.5242 0.5244
Wald tests for linguistic predictors
F 166.5310 55.5259 186.4656 119.5908
d.f. 2.0000 2.0000 2.0000 2.0000
p-value 0.0000 0.0000 0.0000 0.0000
  • Notes: Regression results for predicted outcomes if author is female, with the following dependent variables: rate of citations from papers with a female, male first author and with a female, male last author. These results represent data across all fields. Robust standard errors are shown in parentheses; p-values correspond to two-tailed tests.
@@
Table 5: Regressions predicting citations to sample papers by gender of citing authors
Papers
to two-tailed tests. *p<0.1; **p<0.05; ***p<0.01